In Isosceles trapezoid ABCD, the longer base, AB, is one and one half as long as the shorter bse, CD. The other two sides, AD and BC are both 13cm long.
If the perimeter is 76cm, find the lengths of AB and CD.

Answer by Stitch(43)

Given:
Total perimeter = 76cm = AB + BC + CD + AD
AD & BC = 13cm
AB = 1.5*(CD)
Solution:
Since we know AD & BC are both 13cm long, plug those numbers into the perimeter equation
Place 13cm in for AD & BC
Combine like terms
Subtract 26cm from both sides
Now it is time to substitute the 1.5*(CD) for AB
rewrite the equation
Divide both sides by 2.5

To find AB, plug 20cm in for CD in the equation AB = 1.5*(CD)